# RFDM: Residual Flow Diffusion Model for Efficient Causal Video Editing

### v1

@@ Abstract
Instructional video editing applies edits to an input video using only text prompts, enabling intuitive natural-language control. Despite the rapid progress, most methods still require fixed-length inputs and substantial compute. Meanwhile, autoregressive video generation enables efficient variable-length synthesis, yet remains under-explored for video editing. We introduce a causal, efficient video editing model that edits variable-length videos frame by frame. For efficiency, we start from a 2D image-to-image (I2I) diffusion model and adapt it to video-to-video (V2V) editing by conditioning the edit at time step $t$ on the model’s prediction at t1t-1. To leverage videos’ temporal redundancy, we propose a new I2I diffusion forward process formulation that encourages the model to predict the residual between the target output and the previous prediction. We call this Residual Flow Diffusion Model (RFDM), which focuses the denoising process on changes between consecutive frames. Moreover, we propose a new benchmark that better ranks state-of-the-art methods for editing tasks. Trained on paired video data for global/local style transfer and object removal, RFDM surpasses I2I-based methods and competes with fully spatiotemporal (3D) V2V models, while matching the compute of image models and scaling independently of input video length. More content can be found in RFDM page.
Figure 1: Video style transfer. Editing video style requires understanding of motion and style across frames. Naive Image-to-Image(I2I) models generate inconsistent video frames. Fairy smoothens the result, trading computational cost to achieve lower jittering and inconsistency. RFDM results in the most consistent edit, while being computationally similar to the per-frame baseline (I2I). The video is taken from the DAVIS.
@@ 4.1 Experimental setup
We train RFDM on SD1.5 [33] (RFDM1.5) and SD3.5 (M) [7] (RFDM3.5) as our backbones. For training, we utilize the recently introduced Señorita-2M [56] dataset, which contains 2 million paired videos covering five instructional video editing tasks: global style transfer, local style transfer, object removal, object addition, and object swap. As the dataset lacks predefined train, validation, and test splits, we introduce a fixed random split of 80%, 15%, and 5%, respectively, for consistent training, evaluation, and test. We split the data between train ($80\%$), validation ($5\%$), and test ($15\%$). RFDM is trained on 8 A100 GPUs for 45k steps with a batch size of 8, gradient accumulation of 2, a learning rate of $1\text{e}^{-4}$ using the Adam optimizer [22].
@@ 4.2 Evaluation benchmarks and metrics
Baseline methods.
We compare RFDM against two categories of methods: those built upon fully spatio-temporal (3D) V2V backbones, such as EVE [35], and those based on I2I backbones (2D), such as Fairy [43] and VidToMe [28]. To ensure a fair comparison, for the latter group, which employs the same backbone as ours, we report results using both their original SD1.5 model and our pretrained variant. We use the official repository for VidToMe [28] and reproduce Fairy [43] ourselves, as its implementation is not publicly available.
@@ 4.3 Video editing results
TGVE, TGVE+ and Señorita benchmarks. We compare RFDM against state-of-the-art nobaselines in both editing quality and computational efficiency in Table 1. On TGVE and TGVE+, RFDM3.5 leads in 4 of 6 metrics across the two benchmarks and performs competitively with proprietary V2V models such as EVE, despite relying on a much more efficient I2I backbone. In efficiency, RFDM3.5 matches Fairy’s latency while using $\sim 13\times$ less RAM, and reduces latency by $\sim 4\times$ compared to other baselines. Even the smaller RFDM1.5 remains competitive with I2I state-of-the-art methods such as TokenFlow and RAVE on TGVE/TGVE+, while being $\sim 12$$16\times$ faster and requiring $\sim 4$$5\times$ less RAM. Notably, all RFDM variants achieve the highest CLIPFrame score, suggesting strong temporal consistency, driven by conditioning on $\hat{y}_{t-1}$ (the previous prediction) and predicting a residual flow that updates only changed regions. Compared to EVE, RFDM3.5 attains slightly lower PickScore, which can be explained by EVE’s substantially larger 3D backbone (4.4B vs. 2.5B parameters for SD3.5) and its much larger training set (34M vs. 2M videos in Señorita). Since EVE is closed-source, we cannot directly measure or report its computational cost.
@@ 4.4 Ablations
We conduct ablations on key parameters of RFDM using our smaller variant, SD1.5 backbone, and the Señorita validation set for global style transfer. The number of frames used in the autoregressive loss is fixed to 3 in all ablations, except in the ablation studying this parameter. Our objective is to identify configurations that yield high temporal consistency, low error accumulation, and strong faithfulness to the ground truth. We observe that temporal consistency and error accumulation are discriminative across all ablations except one, for which faithfulness must also be considered. Below, we discuss the impact of each parameter in detail.
@@ 5 Conclusion and Discussion
In this paper, we introduced RFDM, an autoregressive video editing model that adapts an image-to-image backbone through a novel autoregressive conditioning based on residual-flow prediction. Trained on large-scale real-world videos from the Señorita dataset, we extensively ablated across three key aspects: temporal consistency, error accumulation, and faithfulness, and evaluated on both the proposed Señorita benchmark and TGVE+. The results show that causal I2I models like RFDM achieve strong performance and efficiency, pointing to a promising direction for scalable video editing. Limitations: Despite an excellent trade-off between the performance and computational efficiency, RFDM’s drawback lies in its short temporal memory, required in motion-constrained edits such as changing the action of a video. A plausible solution could be deploying a KV caching mechanism, which we leave to future work.
@@ 7.2 Better than ground-truth
We also observe several cases where RFDM produces results that appear more natural than the provided ground truth. Since the Señorita dataset’s ground truth is generated through a multi-stage pipeline combining segmentation, inpainting, and tracking modules, it can occasionally contain artifacts, flickering, or imperfect boundaries. Benefiting from the strong spatial understanding of SD1.5 and the temporal stability introduced by our proposed framework, RFDM often yields smoother and more coherent outputs that correct such imperfections. We encourage the reviewer to refer to the HTML visualization page in the supplementary material for side-by-side video comparisons.

### v2

@@ Abstract
Instructional video editing applies edits to an input video using only text prompts, enabling intuitive natural-language control. Despite the rapid progress, most methods still require fixed-length inputs and substantial compute. Meanwhile, autoregressive video generation enables efficient variable-length synthesis, yet remains under-explored for video editing. We introduce a causal, efficient video editing model that edits variable-length videos frame by frame. For efficiency, we start from a 2D image-to-image (I2I) diffusion model and adapt it to video-to-video (V2V) editing by conditioning the edit at time step $t$ on the model’s prediction at t1t-1. To leverage videos’ temporal redundancy, we propose a new I2I diffusion forward process formulation that encourages the model to predict the residual between the target output and the previous prediction. We call this Residual Flow Diffusion Model (RFDM), which focuses the denoising process on changes between consecutive frames. Moreover, we propose a new benchmark that better ranks state-of-the-art methods by faithfulness for video editing tasks. Trained on paired video data for global/local style transfer and object removal, RFDM surpasses I2I-based methods and competes with fully spatiotemporal (3D) V2V models, while matching the compute of image models and scaling independently of input video length. More content can be found in RFDM page.
Figure 1: Video style transfer. Editing video style requires understanding of motion and style across frames. Naive Image-to-Image(I2I) [34] models generate inconsistent video frames. Fairy [44] smoothens the result, trading computational cost to achieve lower jittering and inconsistency. RFDM results in the most consistent edit, while being computationally similar to the per-frame baseline (I2I). The video is taken from the DAVIS [31].
@@ 4.1 Experimental setup
We train RFDM on SD1.5 [34] (RFDM1.5) and SD3.5 (M) [7] (RFDM3.5) as our backbones. For training, we utilize the recently introduced Señorita-2M [57] dataset, which contains two million paired videos covering five instructional video editing tasks: global style transfer, local style transfer, object removal, object addition, and object swap. As the dataset lacks predefined train, validation, and test splits, we introduce a fixed random split of 80%, 15%, and 5%, respectively, for consistent training, evaluation, and test. We split the data between train ($80\%$), validation ($5\%$), and test ($15\%$). RFDM is trained on 8 A100 GPUs for 45k steps with a batch size of 8, gradient accumulation of 2, a learning rate of $1\text{e}^{-4}$ using the FusedAdam optimizer [22].
@@ 4.2 Evaluation benchmarks and metrics
Baseline methods. We compare RFDM against two categories of methods: those built upon fully spatio-temporal (3D) V2V backbones, such as EVE [36], and those based on I2I backbones (2D), such as Fairy [44] and VidToMe [28]. To ensure a fair comparison, for the latter group, which employs the same backbone as ours, we report results using both their original SD1.5 model and our pretrained variant. We use the official repository for VidToMe [28] and reproduce Fairy [44] ourselves, as its implementation is not publicly available.
@@ 4.3 Video editing results
[…]
@@ 4.4 Ablations
RFDM’s key parameters are ablated using our smaller variant, SD1.5 backbone, and the Señorita validation set for global style transfer. The number of frames used in the autoregressive loss is fixed to 3 in all ablations, except in the ablation studying this parameter. Our objective is to identify configurations that yield high temporal consistency, low error accumulation, and strong faithfulness to the ground truth.
@@ 5 Conclusion and Discussion
In this paper, we introduced RFDM, an autoregressive video editing model that adapts an image-to-image backbone through a novel autoregressive conditioning based on residual-flow prediction. Trained on large-scale real-world videos from the Señorita dataset, we extensively ablated across three key aspects: temporal consistency, error accumulation, and faithfulness, and evaluated on TGVE, TGVE+, and the proposed Señorita benchmark. The results show that causal I2I models like RFDM achieve strong performance and efficiency, pointing to a promising direction for scalable video editing. Limitations: Despite an excellent trade-off between the performance and computational efficiency, RFDM’s drawback lies in its short temporal memory, required in motion-constrained edits such as changing the action of a video. A plausible solution could be deploying a KV caching mechanism, which we leave to future work.
@@ 7.2 Better than ground-truth
We also observe several cases where RFDM produces results that appear more natural than the provided ground truth. Since the Señorita dataset’s ground truth is generated through a multi-stage pipeline combining segmentation, inpainting, and tracking modules, it can occasionally contain artifacts, flickering, or imperfect boundaries. Benefiting from the strong spatial understanding of stable diffusion models and the temporal stability introduced by our proposed framework, RFDM often yields smoother and more coherent outputs that correct such imperfections. We encourage the reviewer to refer to the HTML visualization page in the supplementary material for side-by-side video comparisons.